OpenStudy (anonymous):

The sun radiates about 3.6 x 10 26  joules of energy each second. How much mass does the sun lose each second? @ganeshie8

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

\[E = mc ^{2}\]

OpenStudy (anonymous):

where E is the energy, and m is the mass and C is the speed of light.

OpenStudy (anonymous):

oh thank @IhteshamMalik i need help in this physics .. I'm trying to learn more and I'm doing this practice sheet for school

OpenStudy (anonymous):

hmm good.

OpenStudy (anonymous):

i have some more questions can u help me please!!!

OpenStudy (anonymous):

yup sure.

OpenStudy (anonymous):

@chanlarue the forces are equal because of Force= mass x acceleration. the launcher has a greater mass therefore the acceleration is less. the ball has a greater acceleration and less mass. ball is faster.( m delta t) impulse =mass x velocity = momentum (m delta v) assuming the launcher has more mass, the ball has more momentum

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

Suppose you are on an airplane moving at high speed. If you flip a coin straight up it will land in your lap rather than a great distance behind you. Explain why and include any laws that help prove your point.

OpenStudy (anonymous):

@IhteshamMalik

OpenStudy (anonymous):

Objects in motion stay in motion. It may seem like you aren't in motion, but you're going just as fast as the plane is. It's the same wit the coin, if you hold it in your palm, it isn't going to fly behind you because it's in motion with you. This isn't going to change if you flip it.

OpenStudy (anonymous):

Wait, than why would it fly behind if I put it outside of the plane?" That's all just the wind, if there wasn't any wind (Like inside of the plane) The coin would keep moving. Of course gravity would pull it to the ground though.

OpenStudy (anonymous):

everything in the plane is going at the same speed, including you and the coin!

OpenStudy (anonymous):

It would land in your lap because your not accelerating. one of newtons law states that "An object in motion stays in motion unless acted apon by an outside force."

OpenStudy (anonymous):

soo all of this is the answer?

OpenStudy (anonymous):

No its explanation.... you can check this one.

OpenStudy (anonymous):

it would still land it your lap due to fact the speed of both you and the coin being the same speed and direction and since the coin is up in the air for a fraction of a second, it not slowing down that much Also Newton's first law of motion saying a object ( or body) moving at a constant speed will remain at that speed and direction until some force act on it. Now it you flip the coin and while the coin was in the air accelerate the plane a lot, than the coin would move back toward the back of the plane with F = M* A here a simple experiment you can do Have a friend drive a car ( or van which would be better for this ) and you flip the coin and see where it lands and record the results. do this ten time and graph the results.

OpenStudy (anonymous):

This is the answer

OpenStudy (anonymous):

lol okk

OpenStudy (anonymous):

Derive the time dilation equation found in Section 15.4 of the text. Explain each step of the derivation.

OpenStudy (anonymous):

Time dilation equation is...... \[t=\frac{ t _{o} }{ \sqrt{1-\frac{ V ^{2} }{ C ^{2} }} }\]

OpenStudy (anonymous):

\[t _{o}\] is the proper time.

OpenStudy (anonymous):

t ? is the proper time?

OpenStudy (anonymous):

wat do i use on my keyboard to draw that ?

OpenStudy (anonymous):

down the typing box. click the equation button.. then you can write mathematical expressions

OpenStudy (anonymous):

no i mean on a mac laptop

OpenStudy (anonymous):

You have to install Microsoft offise word...

OpenStudy (anonymous):

ohboy ok

OpenStudy (anonymous):

heres another question

OpenStudy (anonymous):

You sit at the outer rim of a Ferris wheel that rotates at 2 revolutions per minute (RPM). What would your rotational speed be if you were instead clinging to a position halfway from the center to the outer rim?

OpenStudy (anonymous):

Rotational speed will not change. For everyone sitting anywhere on the wheel, it will remain the same, i.e. 2 RPM.

OpenStudy (anonymous):

You sit at the outer rim of a Ferris wheel that rotates at 2 revolutions per minute (RPM). What would your rotational speed be if you were instead clinging to a position halfway from the center to the outer rim?

OpenStudy (anonymous):

Rotational speed will not change. For everyone sitting anywhere on the wheel, it will remain the same, i.e. 2 RPM.

OpenStudy (anonymous):

At the outer edge of a rotating space habitat, 130 m from its center, the rotational acceleration is g. What is the rotational acceleration at a distance of 65 m from the center of the habitat?

OpenStudy (anonymous):

\[\alpha = r \omega ^{2}\]

OpenStudy (anonymous):

and in space a =g = 9.8 meter per second suqare.

OpenStudy (anonymous):

A car traveling at 60 km/h will skid 30 m when its brakes are locked. If the same car is traveling at 180 km/h, what will be its skidding distance?  Show your work.

OpenStudy (anonymous):

\[\alpha = (65)(0.27)^{2}\]

OpenStudy (anonymous):

first you have to wait for the end of the 1st question. then ask next

OpenStudy (anonymous):

ok sorry can we continue this tmm please

OpenStudy (anonymous):

do you get above question's solution?

OpenStudy (anonymous):

find the angular velocity by w^2R = g = 9.81 m/s^2 w^2 = 9.81/130 w = 0.2747 rad/s and at R = 65: a = w^2R = 0.2747^2*65 = 4.905 m/s^2 (= g/2)